### v1

# Continual Multi-Robot Learning from Black-Box Visual Place Recognition Models

Abstract
In the context of visual place recognition (VPR), continual learning (CL) techniques offer significant potential for avoiding catastrophic forgetting when learning new places. However, existing CL methods often focus on knowledge transfer from a known model to a new one, overlooking the existence of unknown black-box models. We explore a novel multi-robot CL approach that enables knowledge transfer from black-box VPR models (teachers), such as those of local robots encountered by traveler robots (students) in unknown environments. Specifically, we introduce Membership Inference Attack, or MIA, the only major privacy attack applicable to black-box models, and leverage it to reconstruct pseudo training sets, which serve as the key knowledge to be exchanged between robots, from black-box VPR models. Furthermore, we aim to overcome the inherently low sampling efficiency of MIA by leveraging insights on place class prediction distribution and un-learned class detection imported from the VPR literature as a prior distribution. We also analyze both the individual effects of these methods and their combined impact. Experimental results demonstrate that our black-box MIA (BB-MIA) approach is remarkably powerful despite its simplicity, significantly enhancing the VPR capability of lower-performing robots through brief communication with other robots. This study contributes to optimizing knowledge sharing between robots in VPR and enhancing autonomy in open-world environments with multi-robot systems that are fault-tolerant and scalable.
I Introduction
Visual place recognition (VPR) enables autonomous robots and self-driving vehicles to classify their place from visual input [1, 2, 3]. While conventional VPR systems rely on supervised learning from direct visual experiences, they face two fundamental limitations: the high cost of collecting training data in new environments and catastrophic forgetting when learning new places. These challenges become particularly acute in long-term autonomous operations where robots must continuously adapt to environmental changes. Recent advancements in continual learning techniques [4]—such as regularization methods, experience replay, and dynamic network expansion—yielded promising results in mitigating catastrophic forgetting and improving continual adaptation performance [5, 6, 7].
This study explores multi-robot continual learning (CL) for visual place recognition (VPR), framing CL as knowledge transfer from an old model to a new one. We extend conventional knowledge transfer between different models of the same robot to knowledge transfer between different models of different robots. This approach offers key advantages, including scalability and fault tolerance. Specifically, a traveling robot can update its knowledge by acquiring information from local robots, avoiding the cost and risk of collecting training data independently. Additionally, sharing VPR knowledge ensures critical information is preserved even if some robots fail. Despite these advantages, multi-robot CL introduces new challenges, such as variations in robot capabilities, communication overhead, and concerns related to privacy and security.
Figure 1: Even simple human-to-human communication or knowledge transfer (KT) allows travelers (students) to avoid getting lost by acquiring VPR knowledge from locals (teachers). In analogy, this study explores a robot-to-robot KT framework. Specifically, we tackle the model inversion problem of reconstructing a (pseudo) training dataset from a black-box teacher with minimal assumptions about the unseen teacher.
One of the key research challenges in extending multi-robot systems is encapsulated in the question: “Can a pseudo-training set be reconstructed from a black-box teacher?” Reconstructing a pseudo-training set from an existing model is a core task in continual learning, as it is essential for integrating new training sets and retraining new models. Many existing approaches rely on the assumption that the existing model is a white box, allowing access to its internal structure or training set. Indeed, in single-robot systems, this assumption often holds, as the existing model belongs to the robot itself. However, in multi-robot systems, especially in open-world multi-robot scenarios where the robots are heterogeneous or unknown robots, making the same assumption is overly optimistic. In fact, reconstructing a training set from a black-box model remains an ongoing area of research in machine learning and is largely an open problem.
In this work, we address the key challenge of reconstructing (pseudo) training samples from black-box teachers using membership inference attacks (MIAs) [8] (Fig. 1). Unlike traditional approaches that analyze the model to reconstruct the training data, MIA generates samples and predicts the likelihood that they were part of the model’s training data. It is the only major privacy attack method that is effective with black-box models. However, this flexibility comes with a downside: low sampling efficiency, particularly when dealing with high-dimensional input data, such as images used in visual place recognition (VPR). The naive approach of randomly generating training sample candidates is highly inefficient. To overcome this limitation, we propose improving the sampling efficiency of MIA by leveraging insights on place class prediction distributions [9] and un-learned class detection [10] from the VPR literature as a prior distribution. We also evaluate the effects of these methods both individually and in combination. Through extensive experiments, we demonstrate that the proposed approach significantly enhances performance in continual learning (CL).
Our key contributions include:
  1. 1.
    This study extends the conventional continual learning (CL) framework from a single-robot to a multi-robot paradigm, facilitating knowledge transfer between robots.
  2. 2.
    It introduces a method for reconstructing pseudo-training samples from a black-box teacher model, tackling challenges in teacher-student VPR knowledge transfer using membership inference attacks (MIA).
  3. 3.
    Experimental results demonstrate that our prior-based approach is remarkably powerful despite its simplicity, significantly enhancing the VPR capability of lower-performing robots through brief communication with other robots.
This study significantly extends our late-breaking paper in [11] and generalizes our previous research on VPR, including continual learning [12], knowledge transfer from teacher to student [9], and knowledge distillation [13], into a novel data-free variant applicable to black-box teachers. As autonomous robots continue to proliferate, universal knowledge transfer protocols and continual domain adaptation will become increasingly important. This research lays the foundation for optimizing robot-to-robot interactions in various VPR scenarios and aims to enhance robotic autonomy in open-world environments.
II Formulation of Multi-Robot CL
This section starts by reviewing the conventional task, which includes the VPR task, supervised learning, and single-robot CL, and builds upon it to formulate multi-robot CL and MIA, which is the focus of our work. Furthermore, as part of the preparation for the evaluation experiments, we introduce a metric for assessing the knowledge transfer cost.
The VPR model is a function $M$ that takes an input image $x$ and outputs the probability distribution $P(y\mid x)$ of the corresponding place class $y$:
$M:x\to P(y\mid x),\quad y\in C.$ (1)
Here, $C$ is a predefined set of place classes, each of which is defined in a real-world region. For instance, in the case of the NCLT public dataset [14] used in our experiments, grid-based partitioning [10] is employed. The workspace is partitioned into $10\times 10$ grid cells in a bird’s-eye coordinate system, with each cell being defined as a place class (Fig. 1). This grid partitioning provides a standard place class; however, as a trade-off, the intra-class variation becomes large, making the classification task more challenging (Fig. 2). One straightforward approach to overcoming the difficulties caused by intra-class variation is optimizing the place definition [15]. This is a fundamental challenge in the VPR community and an ongoing topic of research [3]. While this approach is outside the scope of this paper, we believe it is orthogonal and complementary to the place classification approach presented in this study.
In the traditional supervised learning setup, a robot experiences a set of place classes $C_{0}$, converts this experience into a training set $T_{0}$, and learns a model $M_{0}$ from the training set using supervised learning $L$:
$M_{0}=L(T_{0}).$ (2)
Here, a training set is given in the form:
$T=\{(x_{i},y_{i})\}_{i=1}^{N}.$ (3)
Figure 2: Examples of input images from independent sessions. Each row presents images from different place classes across four sessions, with each column containing four image samples from the corresponding place class. The grid-based partitioning adopted in this paper is the most naive and standard solution to the place class definition issue. However, it results in large intra-class variation, making the VPR task more challenging.
In single-robot CL, when the robot experiences a new class $C_{0}^{+}$ and a new training set $T_{0}^{+}$ arrives, the goal is to update the previous model $M_{0}$ to a new model $M_{1}$. To achieve this, a pseudo training set $\bar{T}_{0}$ is first reconstructed using a model inversion function $I$ (e.g., MIA) as:
$\bar{T}_{0}=I(M_{0}).$ (4)
Here, a pseudo training set is given in the form:
$\bar{T}=\{(x_{i},P(y\mid x_{i}))\}_{i=1}^{N}.$ (5)
It is then combined with the new training set $T_{0}^{+}$ to form:
$\tilde{T}_{0}^{+}=\bar{T}_{0}\cup T_{0}^{+}.$ (6)
Next, the new model is learned using distillation $\bar{L}$ [16] as:
$M_{1}=\bar{L}(\tilde{T}_{0}^{+}).$ (7)
The overall process can be summarized as:
$M_{1}=\bar{L}(I(M_{0})\cup T_{0}^{+}).$ (8)
In multi-robot CL, rather than receiving a new training set $T_{0}^{+}$, a new black-box teacher model $M_{0}^{+}$ arrives. In this case, instead of the generic inversion function $I$, a specialized inversion function $I^{*}$ (i.e., MIA) applicable to the black-box teacher model is required to reconstruct the new training data as:
$\bar{T}_{0}^{+}=I^{*}(M_{0}^{+}).$ (9)
The MIA attempts to reconstruct the training set of a teacher model, which is originally inaccessible, through interactions consisting of queries and responses to the teacher model. The overall process can be expressed as:
$M_{1}=\bar{L}(I(M_{0})\cup I^{*}(M_{0}^{+})).$ (10)
Finally, we discuss the knowledge transfer cost. In multi-robot CL, the interaction between the student robot and the black-box teacher occurs in the form of queries and responses. Specifically, the queries from the student robot are automatically generated by the program code, and the code used in this paper consists of short, fixed-size code snippets of just a few dozen lines. On the other hand, the responses from the teacher are pseudo-training samples. These are represented as long real-valued vectors or tensors, which consume communication costs proportional to the number of pseudo-samples. Based on this background, this paper evaluates the knowledge transfer cost based on the number of pseudo-samples sent from the teacher to the student, assuming that the queries from the student to the teacher do not consume communication costs, or that these costs are negligible.
III Dataset Reconstruction using BB-MIA
A key challenge in applying Membership Inference Attacks (MIA) to black-box models, also known as black-box MIA (BB-MIA), is the high-dimensionality of the sample space. In BB-MIA, the attacker (student) cannot directly access the teacher model’s internal parameters or the distribution of its training data. As a result, generating suitable samples to effectively attack the model’s decision boundaries becomes extremely difficult. To address this issue, we approximate the teacher model using a cascade pipeline consisting of two modules: a pre-trained embedding module and a trainable MIA module. The embedding module transforms high-dimensional input images into lower-dimensional embedding vectors, utilizing a scene graph classifier as an intermediate step, which has been validated for its generalizability and effectiveness across various VPR tasks (e.g., [9]). The MIA module incorporates multiple strategies to efficiently sample from the embedding vectors, enabling high-efficiency MIA sampling. Each module is detailed in the following subsections.
III-A Generic Embedding Model
This pipeline configuration follows our previous research on non-MIA VPR tasks [9]. The pipeline consists of three key stages:
(1) Semantic Segmentation. An image is parsed into image regions by semantic segmentation with DeepLab v3+ [17].
(2) Scene Graph Generation. Image regions are viewed as nodes of a scene graph. Each node is represented as a 189-dimensional one-hot vector containing semantic labels, orientation, and range [15]. Edges are formed by connecting spatially adjacent node pairs. Our scene graph features maintain multimodal regional features, including appearance, semantics, and spatial information, while complementarily combining absolute and relative representations to provide a rich representation. See [18] for details.
(3) Graph Neural Network. Next, the scene graph is passed into a pre-trained graph convolutional network (GCN) classifier, which produces a class-specific probability map of dimension $|C|$ [19]. The classifier’s output is then taken as the final embedding vector.
III-B Black-Box Membership Inference Attacks
The problem here is to reconstruct a pseudo-sample set $\{(x,P(y|x))\}$ that approximates the training data of a black-box teacher, given a teacher that takes an embedding vector as input and outputs a class-specific probability map (CPM). This is an inverse problem of supervised learning.
US (Uniform Sampling) Strategy: The simplest strategy involves sampling $x$ of dimension $|C|$ from a uniform distribution, applying L1 normalization, and inputting it into the teacher model to obtain the output CPM $P(y|x)$.
$x_{i}=\frac{{u}_{i}}{\|{u}_{i}\|_{1}},\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N$ (11)
RR (Reciprocal Rank) Strategy: Although simple, the US strategy is not constrained to approximate the predictive distribution of the classifier model (i.e., scene graph classifier GCN). Therefore, as an alternative to the US strategy, we introduce the reciprocal rank (RR) strategy. As shown in our previous study [9], the output of a classifier model is often approximated by the reciprocal rank feature (RRF), which takes the following form.
$x_{i}=f^{RR}\left(\frac{{u}_{i}}{\|{u}_{i}\|_{1}}\right),\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N$ (12)
where $f^{RR}(x)$ generates a reciprocal rank feature vector by sorting the elements of $x$ in descending order and assigning each element its rank’s reciprocal:
$f^{RR}(x)=\left[\frac{1}{\text{rank}_{1}},\frac{1}{\text{rank}_{2}},\dots,\frac{1}{\text{rank}_{|C|}}\right].$ (13)
Entropy Strategy: Although the RR strategy can approximate the predictive distribution of a generice VPR model, it is not constrained to approximate the predictive distribution of the target teacher model. To overcome this issue, we introduce the Entropy strategy. Entropy is frequently used in the self-localization community to find unseen images for VPR [10]. In contrast, we utilize the entropy measure for the novel application of predicting the place classes that were seen during the training phase of the teacher VPR model. The Entropy strategy selects high-quality samples based on the assumption that samples $(x,P(y|x))$ with low entropy of $P(y|x)$ are more likely to be members of the teacher model’s training set. Specifically, we generate an excessive number of samples using the RR strategy and select the top $N$ samples with the lowest entropy score $H$.
$H(x)=-\sum_{i=1}^{|C|}P(y|x_{i},f)\log P(y|x_{i},f).$ (14)
Replay/Prior Strategy: Leaving the category of BB-MIA methods, we also consider the conventional replay-based sampler for benchmarking purposes, which contradicts the assumption of a black-box teacher [20]. This strategy assumes that the teacher is not a complete black-box and retains a subset of the training dataset bundled with the teacher model, and that the student has access to this subset. Such an assumption is relevant in multi-robot systems with superior communication and memory capacities, where the communication channel between the teacher and student is sufficiently wide, and the teacher model’s memory capacity is large enough. As a variant of the standard replay strategy in continual learning, we also introduce a novel strategy called the Prior strategy. Instead of assuming that the teacher model retains its training samples, as in the replay strategy, the Prior strategy assumes that the student model is bundled with its own training samples. These available training samples are then used as additional samples for querying the teacher. Such an assumption is relevant in scenarios where only the student, rather than the teacher, is cooperative in continual learning and where the communication bandwidth is sufficiently large to transmit the training samples.
Mixup Strategy: In addition, we introduce the Mixup strategy, which combines the advantages of the replay strategy and other (RR or Entropy) strategies. This strategy assumes that only a small subset of the training set maintained by the student/teacher model, with a size of $R$ (e.g., $R=1$), is retained. This assumption is relevant when the communication and memory capabilities of a multi-robot system are not as abundant as those envisioned by the Replay strategy, but are still not entirely unavailable. It generates the required $N$ samples by combining the retained $R$ samples with $(N-R)$ samples from the set generated by the RR/Entropy strategy:
${x}_{i}\sim\begin{cases}T^{Replay}&\text{if }i=1,\dots,R,\\ T^{RR/Entropy}&\text{if }i=R+1,\dots,N.\end{cases}$ (15)
This strategy is clearly inapplicable to a black-box teacher and should therefore be used only as an oracle baseline for benchmarking purposes.
IV Experimental Evaluation
We evaluate the performance of the proposed method in a typical multi-robot continual learning (CL) scenario, where a traveler robot (the student) encounters three teachers in succession and receives knowledge transfer via wireless communication. However, there is also the risk of forgetting previously learned place classes. One of our goals is to investigate the trade-off between knowledge acquisition and forgetting.
IV-A Experimental Setup
We used the NCLT dataset [14] in our experiments. The NCLT dataset provides sensor data obtained from a Segway robot navigating across multiple sessions spanning across multiple seasons on the North Campus of the University of Michigan. Specifically, in our VPR tasks, we use images from the robot’s onboard camera as sensor input, annotated with ground-truth viewpoint GPS data. We follow a common protocol for using the NCLT dataset, collected by a single robot, in multi-robot scenarios, where different robots are paired with different sessions [21].
Our evaluation protocol:
  1. 1.
    Sequential teacher interaction (up to 3 teachers)
  2. 2.
    27 NCLT dataset sessions
    • 1 test session: “2012/08/04”
    • 1 training session for embedding model: “2012/04/29”
    • 25 sessions for student/teacher VPR model training
  3. 3.
    Teacher/Student VPR models
    • MLP with a hidden layer of 4,096 dimensions
  4. 4.
    6 distinct knowledge transfer scenarios
  5. 5.
    Performance metrics:
    • Top-1 VPR accuracy measured at:
      • After the student’s supervised learning
      • After knowledge transfer from the 1st teacher
      • After knowledge transfer from the 2nd teacher
      • After knowledge transfer from the 3rd teacher
    • KT cost:
      • The number of samples $N$ used for KT
    • Knowledge retention
      • Avoidance of catastrophic forgetting
    • Computational efficiency
Figure 3: Example of experimental setup. In each scenario, at stage $i=0$, the student robot trains the VPR model via supervised learning using the data from the place classes it has experienced (blue boxes). In subsequent stages $i=1,2,3$, every time the student encounters a new teacher, it retrains the VPR model via knowledge transfer using the data of the place classes the teacher has experienced (yellow boxes).
Several detailed setup configurations are shown below: (1) The student robot and teacher robots have experienced $K$ place classes and have trained a place classifier in a supervised manner using the full training set for these experienced place classes from the corresponding session. Unless otherwise specified, following the standard continual learning protocol, once the training is completed, the training set is not retained and is discarded for all BB-MIA strategies except Replay, Prior, and Mixup. (2) Session IDs from 0 to 24 were assigned to 25 different sessions, in order of the navigation dates, starting from the earliest. (3) To evaluate performance across different student-teacher combinations, we introduce different scenarios for $j=0,1,\dots,5$. In the $j$-th scenario, the student and the three teachers are distinguished by model ID $i$. Specifically, $i=0$ represents the student model, while $i=1,2,3$ correspond to the $i$-th teacher. Each $i$-th model experiences $K$ random place classes during the $((6i+j)\mod 25)$-th session among the 25 sessions. By default, $K=10$. (4) For simplicity, the number of samples appearing in this section will refer to the number of samples per place class, rather than the total number across all place classes. (5) When the student encounters a teacher, (pseudo) training samples are reconstructed from the teacher model for place classes known to the teacher, while for place classes known only to the student, training samples are reconstructed from the student model. (6) In the Replay, Prior and Mixup strategy, the number of mixed replay samples $R$ per class was set to $R=1$ by default. (7) RRF vectors were approximated using a sparse $k$-hot RRF ($k=10$) following the method in [19].
The computational cost of this experiment was lightweight, with training times of tens of seconds for MLP models and about 25 milliseconds per sample for question generation. Additionally, the KT cost of transferring a 100-dimensional $k$-hot RRF sample was less than 128 bits.
IV-B Results and Discussions
First, we evaluated the basic performance of the proposed method. We examined the performance at the initial stage when the student was trained with supervised learning, as well as at later stages when the student continued learning through knowledge transfer from new teachers.
At the initial stage, when the student had not yet encountered any teachers, its performance was predictably poor. This was because test sessions included questions from unseen place classes, making it evident that even the most capable student would struggle to provide accurate answers for such test samples.
Figure 4: Top-1 accuracy vs. KT cost ($N$).
Figure 4 shows that the performance improved with an increasing number of samples across all five strategies, but differences among them were observed:
Replay Strategy: When the number of samples was sufficiently large, this strategy remained unaffected by catastrophic forgetting [22] and achieved the highest performance across all experiments. However, it comes at the cost of not adhering to the principles of CL. The Prior strategy, a variant of the Replay strategy, which queries the teacher using the student’s training samples, did not perform well. This may be due to the low probability of the student’s training sample coinciding with the teacher’s training sample by chance.
US Strategy: This strategy exhibited the lowest performance in all experiments, likely due to the non-uniform sample distribution in the embedding vector space.
RR Strategy: Despite being simple, this strategy surprisingly demonstrated high performance. This result suggests that the RRF distribution serves as a good approximation of the prediction distribution.
Entropy Strategy: While being still simple, it performed comparably or better than the RR strategy, especially excelling at generating elite samples when $N$ was small.
Mixup Strategy: This strategy balances generalization and cost-effectiveness.
As mentioned earlier, both Replay and Mixup do not satisfy the assumption of a black-box teacher and are used solely for benchmarking purposes. Although Mixup requires retaining a small number ($R$) of training samples, its knowledge transfer cost is significantly lower than that of Replay, mitigating catastrophic forgetting while achieving performance close to that of the Replay strategy.
V Conclusion and Future Work
We formulated the problem of continual multi-robot learning for visual place recognition (VPR) as teacher-student knowledge transfer and proposed exploiting membership inference attacks (MIA), the only major privacy attack capable of handling black-box teachers. To enable the use of MIA for knowledge transfer, we focused on the key challenge of its low sampling efficiency, which acts as a bottleneck. We presented a method that leverages the student model’s prior knowledge to achieve practical sampling efficiency. Through extensive experiments, we investigated the relationship between VPR performance, sampling efficiency, and computational efficiency, demonstrating the significant effectiveness of our proposed approach.
In this study, we focused on a fundamental investigation, addressing a near-worst-case scenario where all teachers are black boxes and all robots have low performance. For practical deployment, future research must urgently explore heterogeneous multi-robot systems that include white-box teachers and robots equipped with specialized continual learning capabilities. Furthermore, while this study adopted a highly naive grid-based partitioning for place definition, prior research suggests that a straightforward extension using more advanced place definition methodologies could dramatically improve visual place recognition performance. Another promising research direction is extending beyond single-image-based visual place recognition to robust self-localization using image sequences and particle filtering, which offers a guaranteed performance boost.